OpenStudy (anonymous):

solve the compound inequality 7>-2x+6 or 10 less than or equal to -2x+3

OpenStudy (a_clan):

eqn 1.\[7>-2x+6\] x>-1/2 .....(1) eqn 2. \[10 \le -2x+3\] \[x \le-7/2\].....(2) \[x=(-\infty,-7/2]\cup(-1/2,\infty)\]

OpenStudy (anonymous):

\[-{1\over2}<x \le-{7\over2}\]